PITMinutes 05-23-2016

From CASA Wiki
Jump to: navigation, search


R. Bachimanchi, M. Drury, A. Freyberger, C. Hovater, A. Hutton, G. Krafft, A. Kimber, C. Mounts, R. Nelson, T. Plawski, C. Reece

Agenda and Notes

1. Operations Status and Summary

B receiving beam again!

Cryo hiccups resolved with work

Heater testing in R100

Decarad measurements with all C100s

Microphonic damping feet on 0L04: didn't fit keep trying

2. FE/Particulate Meeting Report (Geng)

Follow on activity and future planning

Studies of warm girders next to the removed 1L12 cryomodule.

Day time scales for looking. After Prad studies In July!

Dummy girder? Gives us maybe 2 weeks to study removed girder.

Replace ion pumps? Clean interior of both sides?

3. Capture degradation rate?

Gradient calibrations have changed

Accounting for this, there hasn't been serious degradation?

monitoring going forward. Data archived.

Look at archiver data and compare?

Other ideas?

Action Items